The leachate pond consists of tropic acid at a concentration of 3.5 x 10-4 M. If the combustion reaction in d(i) releases -1215kJ, calculate the temperature change of the pond assuming that the pond has a volume of 15000L, density of 1080g/L and its specific heat is the same as pure water.          


1
Expert's answer
2021-11-29T04:37:30-0500

Mass = 1500×1080=1620000g=1620kg


The specific heat capacity of water is 4,200 Joules per kilogram per degree Celsius (J/kg°C). 


Hear {Q} = -1215kJ


Q= mc∆T

"∆T=\frac{1215}{1620×4200}=0.00018°c"